OpenStudy (jackiem272):

dy/dt= y-t^2 ..how would you get y(t)?

OpenStudy (fortytherapper):

Are you wanting to integrate?

OpenStudy (jackiem272):

the problem asks me to use the graph of the solution y(t), and dy/dx is given

OpenStudy (jackiem272):

i think i have to integrate

OpenStudy (fortytherapper):

Think you can upload the picture here some way?

OpenStudy (jackiem272):

http://math.sci.ccny.cuny.edu/document/show/2156

OpenStudy (jackiem272):

page 47, #18c

geerky42 (geerky42):

Yeah just integral both sides with respect to t

OpenStudy (fortytherapper):

Yeah, let's first integrate to get back to the original \[\int\limits_{}^{}y-t^2\] So I guess \[\int\limits_{}^{}y - \int\limits_{}^{}t^2\]

OpenStudy (jackiem272):

i was wondering if i could do that...but for some reason i felt it was wrong

geerky42 (geerky42):

Then you will obtain equation with constant. You were given \(y(0)=1\) so plug in y = 1 and t = 0 then solve for c

OpenStudy (freckles):

if you want to solve for y you will have to find the integrating factor

OpenStudy (freckles):

\[y'-y=-t^2 \\ y'+\color{red}{(-1)}y=-t^2 \\ \text{ the integrating factor is } e^{\int\limits_0^t -1 dx}=e^{-t}\] \[e^{-t} y'+(-1)e^{-t}y=e^{-t}t^2 \\ (e^{-t}y)'=e^{-t}t^2\] now you can integrate both sides

OpenStudy (jackiem272):

is that how i will find the original equation?

OpenStudy (solomonzelman):

this is a first order, linear (non-separable) differential equation. that is how it is classified, but to understand this, and understand it completely, all you need to know is: \(\color{#000000}{ \small \displaystyle [1]\quad\quad \frac{d}{dx}e^{f(x)}=e^{f(x)}\times f'(x) }\) (via Chain Rule) \(\color{#000000}{ \small \displaystyle [2]\quad\quad \frac{d}{dx}\left[f(x)g(x)\right]=f'(x)g(x)+f(x)g'(x) }\) (definition of Product Rule)

OpenStudy (solomonzelman):

If you know these two things, then I can proceed to explaining how to solve for \(y(t)\) in such problems. So, do you, or not?

OpenStudy (jackiem272):

ok, yes i understand it